OpenStudy (anonymous):

Multiplying and Dividing Mixed Numbers What is 5*1 3/5 ? A.5 2/5 B.5 /38 C.8 D. 15

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@MathiznotEZ

TheSmartOne (thesmartone):

Hi there! :)

OpenStudy (anonymous):

hi

TheSmartOne (thesmartone):

Let's make 1 3/5 in to an improper fraction :)

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

Do you know how to do that? :)

OpenStudy (anonymous):

you divide right?

TheSmartOne (thesmartone):

nope

OpenStudy (anonymous):

oh im lost then

TheSmartOne (thesmartone):

for example to make 2 3/4 into improper fraction, I would do this: \(\sf\Large \frac{(2\times 4) + 3}{4} = \frac{8+3}{4} = \frac{11}{4}\)

OpenStudy (anonymous):

i got 8/5

OpenStudy (anonymous):

\[\frac{ 5*1+3 }{ 5 }=\frac{ 8 }{ 5 }\] thats what i got

TheSmartOne (thesmartone):

good job! :D

TheSmartOne (thesmartone):

and now we multiply :)

OpenStudy (anonymous):

thank you

TheSmartOne (thesmartone):

\(\sf\Large 5\times \frac{8}{5} = \frac{5}{1} \times \frac{8}{5} =?\)

OpenStudy (anonymous):

= 33/5 or = 6 3/5 right?

TheSmartOne (thesmartone):

mhmm, how did you get 33?

OpenStudy (anonymous):

oh woops i did it wrong lol

OpenStudy (anonymous):

sorry

TheSmartOne (thesmartone):

Remember, the rule for multiplying fractions: \(\sf\Large \frac{a}{b} \times \frac{c}{d} = \frac{ac}{bd}\)

OpenStudy (anonymous):

8/1 or 8?

OpenStudy (anonymous):

its c?

TheSmartOne (thesmartone):

Anything divided by 1 is the same thing. So 8 is correct :)

TheSmartOne (thesmartone):

c is correct :D

OpenStudy (anonymous):

yay thank you

TheSmartOne (thesmartone):

Anytime! And great work! :D

OpenStudy (anonymous):

thank you you did good too

OpenStudy (anonymous):

no not good awesome! :)

TheSmartOne (thesmartone):

:)